A school sports team contains 68 students. 33 do field events, 40 do track events, 23 do swimming, 14 do both field and track events, 8 do both swimming and field events. If 15 students do field events only and 10 do both swimming and track events, how many students do
a) swimming only,
b) track events only,
c) all three events?

Answer by ikleyn(52778)

.
A school sports team contains 68 students.
33 do field events, 40 do track events, 23 do swimming,
14 do both field and track events, 8 do both swimming and field events.
If 15 students do field events only and 10 do both swimming and track events, how many students do
a) swimming only,
b) track events only,
c) all three events?
~~~~~~~~~~~~~~~~~~~~~~

We are given a universal set of 68 students and 3 its basic subsets 

    F  of 33 students

    T  of 40 students

    S  of 23 students.


We also are given three in-pair intersections

    FT of 14 students

    FS of  8 students

    TS of 10 students.


We also are given that the set

    F_only contains 15 students.


Let x be the number of students that do all 3 activities FTS.


Then, due to the Inclusion-Exclusion principle, we have this equation

    68 - n(F) + n(T) + n(S) - n(FT) - n(FS) - n(TS) + x,

or

    68 = 33 + 40 + 23 - 14 - 8 - 10 + x.


From it, we finf

    x = 68 - 33 - 40 - 23 + 14 + 8 + 10 = 4.


So, the ANSWER to (c) is 4 students do all 3 activities.


Now for (a), "swimming only" has  S_only = S - SF - TS + FTS = 23 -  8 - 10 + 4 =  9  students.   ANSWER


Now for (b), "track only"    has  T_only = T - FT - TS + FTS = 40 - 14 - 10 + 4 = 20 students.    ANSWER


At this point, the problem is solved in full. 
All questions are answered.

The solution is complete.